# Comparative Analysis of LLM Abliteration Methods: A Cross-Architecture Evaluation

### v1

@@ Abstract
Safety alignment mechanisms in large language models prevent responses to harmful queries through learned refusal behavior, yet these same mechanisms impede legitimate research applications including cognitive modeling, adversarial testing, and security analysis. While abliteration techniques enable surgical removal of refusal representations through directional orthogonalization, the relative effectiveness of available implementations remains uncharacterized. This study evaluates four abliteration tools (Heretic, DECCP, ErisForge, FailSpy) across sixteen instruction-tuned models (7B–14B parameters), reporting tool compatibility on all 16 models and quantitative metrics on subsets dictated by tool support (Heretic: KL divergence + refusal rate on eight models; benchmark preservation on three models with baselines plus Zephyr-7B-beta). Single-pass methods demonstrated superior capability preservation on the benchmarked subset (avg GSM8K $\Delta$ across three models: ErisForge 0.28-0.28 $p$; DECCP 0.13-0.13 $p$), while Bayesian-optimized abliteration produced variable distribution shift (KL divergence: 0.043–1.646) with model-dependent capability impact. These findings provide researchers with evidence-based selection criteria for abliteration tool deployment across diverse model architectures. The principal finding indicates that mathematical reasoning capabilities exhibit the highest sensitivity to abliteration interventions, with GSM8K change ranging from +1.51+1.51 $p$ to 18.81-18.81 $p$ ($26.5\%$ relative) depending on tool selection and model architecture.
@@ 3.3 Experimental Protocol
Statistical Analysis: Benchmark scores are reported as mean accuracy percentages. For cross-tool comparisons, percentage point change from baseline ($\Delta=\text{Abliterated}-\text{Baseline}$) is computed to isolate abliteration effects. Effect magnitudes are interpreted as: minimal ($|\Delta|<1$ $p$), moderate ($1\leq|\Delta|<5$ $p$), and substantial ($|\Delta|\geq 5$ $p$). KL divergence values are computed using first-token probability distributions averaged across 100 harmless prompts. Benchmark uncertainty is summarized via standard errors reported by lm-evaluation-harness; Appendix reports 95% confidence intervals computed as score $\pm 1.96\times\text{stderr}$. Benchmark uncertainty: lm-evaluation-harness reports standard errors for benchmark metrics. In the benchmark runs analyzed here, standard errors were 0.38–0.40 $p$ for MMLU, 1.23–1.38 $p$ for GSM8K, and 0.37–0.42 $p$ for HellaSwag, providing a rough scale for interpreting small differences.
@@ 4.2 Benchmark Preservation Results
Table 3 presents capability preservation using MMLU, GSM8K, and HellaSwag benchmarks. Given the reported benchmark standard errors (Methods), small differences ($\sim$1 $p$) on MMLU and HellaSwag and modest differences on GSM8K should be interpreted cautiously. In contrast, the Yi-1.5-9B GSM8K drop under Heretic (18.81-18.81 $p$) is an order of magnitude larger than the GSM8K standard error, indicating a substantial capability impact rather than sampling noise.
@@
Benchmarks from lm-evaluation-harness; typical standard errors are $\sim$0.4 $p$ (MMLU), $\sim$1.3 $p$ (GSM8K), $\sim$0.4 $p$ (HellaSwag).
@@ 5.1 Tool Selection Considerations
Capability preservation priority: ErisForge and DECCP showed the lowest benchmark degradation (avg GSM8K $\Delta$: 0.28-0.28 $p$ and 0.13-0.13 $p$ respectively).
@@ 5.3 Limitations
Statistical Rigor: Due to computational constraints, the primary results reflect single experimental runs rather than multiple independent trials with variance estimation. While Heretic’s Bayesian optimization internally evaluates multiple configurations (50 trials), each final abliterated model represents a single optimization outcome. This design choice was driven by practical considerations: full benchmark evaluation of a single 7B model requires approximately 4 to 6 GPU-hours, making triplicate runs across 16 models and 4 tools prohibitively expensive (>>1,500 GPU-hours). To mitigate this limitation, deterministic execution was ensured (fixed seeds, deterministic CUDA where possible) and complete reproducibility artifacts are provided enabling independent replication. Future work with dedicated compute resources should incorporate proper bootstrap confidence intervals and statistical significance testing across multiple runs. Notably, the observed effect sizes for tool comparisons (e.g., Yi-1.5-9B GSM8K: Heretic 52.08% vs DECCP 72.40%, $\Delta=20.32$ $p$) substantially exceed benchmark uncertainty (GSM8K 95% CI half-width $\sim$2.4–2.7 $p$; Appendix), suggesting robust underlying differences despite single-run methodology.
@@ 6 Conclusion
1. Capability preservation varied significantly by tool. In the benchmarked subset, ErisForge exhibited the lowest average benchmark degradation (MMLU: 0.12-0.12 $p$, GSM8K: 0.28-0.28 $p$, HellaSwag: 0.08-0.08 $p$), followed by DECCP (GSM8K: 0.13-0.13 $p$). Heretic demonstrated model-specific variance, with minimal impact on Mistral-7B (0.12-0.12 $p$ avg) but substantial GSM8K degradation on Yi-1.5-9B (18.81-18.81 $p$; $26.5\%$ relative).
2. Mathematical reasoning showed highest sensitivity to abliteration. GSM8K scores exhibited the largest variance across tools and models (range: +1.51+1.51 $p$ to 18.81-18.81 $p$), suggesting potential overlap between mathematical reasoning circuits and refusal-related representations.
@@
5. Model resilience varied across architectures. Mistral-7B exhibited consistent stability across all tools (avg $\Delta$: 0.12-0.12 to 0.61-0.61 $p$), while Yi-1.5-9B showed high sensitivity to Heretic’s kernel-weighted approach but improved GSM8K performance (+1.51+1.51 $p$) with DECCP.

### v2

@@ Abstract
Safety alignment mechanisms in large language models prevent responses to harmful queries through learned refusal behavior, yet these same mechanisms impede legitimate research applications including cognitive modeling, adversarial testing, and security analysis. While abliteration techniques enable surgical removal of refusal representations through directional orthogonalization, the relative effectiveness of available implementations remains uncharacterized. This study evaluates four abliteration tools (Heretic, DECCP, ErisForge, FailSpy) across sixteen instruction-tuned models (7B–14B parameters), reporting tool compatibility on all 16 models and quantitative metrics on subsets dictated by tool support (Heretic: KL divergence + refusal rate on eight models; benchmark preservation on three models with baselines plus Zephyr-7B-beta). Single-pass methods demonstrated superior capability preservation on the benchmarked subset (avg GSM8K $\Delta$ across three models: ErisForge 0.28-0.28 pp; DECCP 0.13-0.13 pp), while Bayesian-optimized abliteration produced variable distribution shift (KL divergence: 0.043–1.646) with model-dependent capability impact. These findings provide researchers with evidence-based selection criteria for abliteration tool deployment across diverse model architectures. The principal finding indicates that mathematical reasoning capabilities exhibit the highest sensitivity to abliteration interventions, with GSM8K change ranging from +1.51+1.51 pp to 18.81-18.81 pp ($26.5\%$ relative) depending on tool selection and model architecture.
@@ 3.3 Experimental Protocol
Statistical Analysis: Benchmark scores are reported as mean accuracy percentages. For cross-tool comparisons, percentage point change from baseline ($\Delta=\text{Abliterated}-\text{Baseline}$) is computed to isolate abliteration effects. Effect magnitudes are interpreted as: minimal ($|\Delta|<1$ pp), moderate ($1\leq|\Delta|<5$ pp), and substantial ($|\Delta|\geq 5$ pp). KL divergence values are computed using first-token probability distributions averaged across 100 harmless prompts. Benchmark uncertainty is summarized via standard errors reported by lm-evaluation-harness; Appendix reports 95% confidence intervals computed as score $\pm 1.96\times\text{stderr}$. Benchmark uncertainty: lm-evaluation-harness reports standard errors for benchmark metrics. In the benchmark runs analyzed here, standard errors were 0.38–0.40 pp for MMLU, 1.23–1.38 pp for GSM8K, and 0.37–0.42 pp for HellaSwag, providing a rough scale for interpreting small differences.
@@ 4.2 Benchmark Preservation Results
Table 3 presents capability preservation using MMLU, GSM8K, and HellaSwag benchmarks. Given the reported benchmark standard errors (Methods), small differences ($\sim$1 pp) on MMLU and HellaSwag and modest differences on GSM8K should be interpreted cautiously. In contrast, the Yi-1.5-9B GSM8K drop under Heretic (18.81-18.81 pp) is an order of magnitude larger than the GSM8K standard error, indicating a substantial capability impact rather than sampling noise.
@@
Benchmarks from lm-evaluation-harness; typical standard errors are $\sim$0.4 pp (MMLU), $\sim$1.3 pp (GSM8K), $\sim$0.4 pp (HellaSwag).
@@ 5.1 Tool Selection Considerations
Capability preservation priority: ErisForge and DECCP showed the lowest benchmark degradation (avg GSM8K $\Delta$: 0.28-0.28 pp and 0.13-0.13 pp respectively).
@@ 5.3 Limitations
Statistical Rigor: Due to computational constraints, the primary results reflect single experimental runs rather than multiple independent trials with variance estimation. While Heretic’s Bayesian optimization internally evaluates multiple configurations (50 trials), each final abliterated model represents a single optimization outcome. This design choice was driven by practical considerations: full benchmark evaluation of a single 7B model requires approximately 4 to 6 GPU-hours, making triplicate runs across 16 models and 4 tools prohibitively expensive (>>1,500 GPU-hours). To mitigate this limitation, deterministic execution was ensured (fixed seeds, deterministic CUDA where possible) and complete reproducibility artifacts are provided enabling independent replication. Future work with dedicated compute resources should incorporate proper bootstrap confidence intervals and statistical significance testing across multiple runs. Notably, the observed effect sizes for tool comparisons (e.g., Yi-1.5-9B GSM8K: Heretic 52.08% vs DECCP 72.40%, $\Delta=20.32$ pp) substantially exceed benchmark uncertainty (GSM8K 95% CI half-width $\sim$2.4–2.7 pp; Appendix), suggesting robust underlying differences despite single-run methodology.
@@ 6 Conclusion
1. Capability preservation varied significantly by tool. In the benchmarked subset, ErisForge exhibited the lowest average benchmark degradation (MMLU: 0.12-0.12 pp, GSM8K: 0.28-0.28 pp, HellaSwag: 0.08-0.08 pp), followed by DECCP (GSM8K: 0.13-0.13 pp). Heretic demonstrated model-specific variance, with minimal impact on Mistral-7B (0.12-0.12 pp avg) but substantial GSM8K degradation on Yi-1.5-9B (18.81-18.81 $p$; $26.5\%$ relative).
2. Mathematical reasoning showed highest sensitivity to abliteration. GSM8K scores exhibited the largest variance across tools and models (range: +1.51+1.51 pp to 18.81-18.81 pp), suggesting potential overlap between mathematical reasoning circuits and refusal-related representations.
@@
5. Model resilience varied across architectures. Mistral-7B exhibited consistent stability across all tools (avg $\Delta$: 0.12-0.12 to 0.61-0.61 pp), while Yi-1.5-9B showed high sensitivity to Heretic’s kernel-weighted approach but improved GSM8K performance (+1.51+1.51 pp) with DECCP.
@@
Future Directions. Several promising extensions merit investigation. Recent contributions to the llm-abliteration repository [5] have extended support to Mixture-of-Experts (MoE) architectures, enabling abliteration of models such as Mixtral and DeepSeek-MoE that were previously incompatible with transformer-focused tools. Additionally, the introduction of an “invert” option enables directional addition rather than removal, potentially amplifying specific behavioral traits—a capability that opens novel applications in steering model personality, enhancing helpfulness, or studying the symmetric effects of representation engineering. Systematic evaluation of these extensions across MoE architectures and bidirectional modification scenarios represents a natural continuation of this work.